# DISPATCH: Distilling Selective Patches for Speech Enhancement

### v1

@@ 1 Introduction
In recent years, deep neural networks have demonstrated strong performance in speech enhancement [1] [2] [3]. Deeper and more sophisticated network architectures have driven these gains, but greater computational power and memory are required. Such computational demands pose significant challenges for deployment in practical applications, particularly in resource-constrained settings such as on-device AI, where hardware limitations, power consumption, and monetary costs are critical concerns. Consequently, model compression has become a crucial research direction for developing computationally efficient speech enhancement models. Several compression strategies have been explored, including pruning [4], which removes redundant weights; quantization [5], which reduces the numerical precision of model parameters; and Knowledge Distillation (KD) [6] [7] [8]. This paper focuses on KD, a teacher-student paradigm in which a compact student model is trained to mimic a larger teacher model, thereby inheriting the teacher’s knowledge.
@@
KD has been actively explored in the speech domain, such as speaker verification [9] [10], speech separation [11] [12] , and speech enhancement [13] [14] [15] [16] [17]. A prominent line of work is response-based KD, which trains a student to mimic the teacher’s output [9] [11] [12] [13] [17]. This output can be diverse, ranging from logits or posterior probabilities to time-frequency output. Some approaches [10] [14] [15] combine response-based and feature-based KD to leverage complementary information at different levels. These approaches simultaneously employ both strategies, where the student learns not only to match the teacher’s final output but also to align intermediate feature representations.
@@
In speech enhancement, distilling time–frequency output has become a prominent response-based KD strategy because this approach can effectively capture the rich time-frequency structure of speech. This approach has given rise to notable frequency-dependent methods, such as Subband-KD [13] and Dynamic Frequency-Adaptive Knowledge Distillation (DFKD) [17]. DFKD addresses the limitation of Subband-KD arising from static band partitioning by introducing a frequency adapter that dynamically partitions the spectrogram into low- and high-frequency regions and optimizes each region with distinct losses. This dynamic frequency-dependent approach allows DFKD to significantly outperform Subband-KD and ABC-KD [15], a hybrid of response- and feature-based distillation. Although DFKD has shown strong performance, it remains limited by relying on the entire output of the teacher, which can contain low‑quality time–frequency regions.
@@
KD rests on the assumption that the teacher’s knowledge is superior to the student’s. Under this premise, the traditional method compels a student to replicate the teacher’s output indiscriminately. However, this strategy has two critical limitations. The first limitation concerns the student’s existing proficiency. Distilling knowledge in regions where the student has already converged yields minimal benefit. The second challenge stems from the inherent imperfections of the teacher model, which may result from overfitting to the training dataset. Therefore, forcing the student to learn from flawed regions in the teacher’s output is detrimental, as it leads the student to replicate the teacher’s weakness instead of its strengths. As a result, indiscriminate distillation not only diverts the student’s focus from the regions where the teacher’s guidance is most needed but also, more importantly, can cause the student to imitate the low-quality output, leading to a degradation in performance.
@@
Fig. 1: Overview of DISPatch framework: DISPatch framework separates spectrograms into patches and selectively applies KD loss to those exhibiting a high Knowledge Gap, where the teacher’s error is low (L) and the student’s is high (H). Patch size: $2C\times N\times 1$ .
@@
A recent work on large language models has demonstrated that their accelerated convergence and improved performance can be achieved through the selective training strategies [18]. This approach involves training the model exclusively on useful tokens, which are identified by a reference model trained on a high-quality dataset. Inspired by this principle, we propose Distilling Selective Patches (DISPatch). This strategy distills only the most useful spectrogram patches, leveraging the notion of selectivity in KD and mitigating the limitations discussed above. The overall architecture of DISPatch is illustrated in Fig. 1. We introduce a Knowledge Gap Score to enable a more targeted transfer of the teacher’s knowledge by identifying and distilling only the most valuable patches, rather than imposing the KD loss across the entire spectrogram. DISPatch explicitly ignores regions where the teacher is unreliable or where the student is already proficient. Moreover, recognizing that speech characteristics vary across frequencies, we introduce Multi-Scale Selective Patches (MSSP). This is a frequency-dependent strategy that adapts the distinct patch sizes to different frequency regions for more tailored distillation. We show consistent gains by applying DISPatch to response-based KD, and further improvements when integrating DISPatch and MSSP into DFKD.
@@ 2.1 Not All Spectrogram Regions Are Equal
Fig. 2: The average error of patches by category. (a) and (b) show the average error for different patch types at the same speech sample for the teacher and student, respectively.
@@
Unlike the common belief that the teacher’s knowledge is superior to the student’s, this analysis reveals heterogeneous behaviors across individual patches. Fig. 2, which presents average error for different types of patches in the same speech sample, shows that the teacher is not flawless; as seen in Fig. 2 (a), it contains low-quality regions such as Challenging (0.8%) and Exploding (3.5%) patches. Conversely, Fig. 2 (b) indicates that a substantial portion of the student’s patches are already high-quality, being classified as Converged (70.4%) patches. This analysis highlights the limitations of indiscriminately applying KD loss across the entire teacher’s output. This conventional approach presents two significant problems. First, it forces the student to imitate the teacher’s low-quality output (Challenging or Exploding). Second, it diverts the student’s focus to regions where the student already excels (Converged), rather than where improvement is needed (Challenging or Exploding).
@@ 2.2 DISPatch: Distilling Selective Patches
2.3 Multi-Scale Selective Patches
@@ 3.2 Implementation Details
We evaluate DISPatch using ConvTasNet (Deep w/ PReLU) [19] and DCCRN-CL [2]. For the implementation of ConvTasNet (Deep w/ PReLU), the teacher (9.71M) and student (1.75M) are configured with hyperparameters [N, L, B, H, Sc, P, X, R] set to [512, 16, 128, 512, 128, 3, 8, 3] and [128, 40, 128, 256, 128, 3, 7, 2], respectively. The DCCRN-CL teacher (9.22M) and student (2.24M) are configured with encoder channels of [64, 128, 256, 256, 512, 512] and [32, 64, 128, 128, 256, 256], respectively. Additionally, the teacher is configured with two complex LSTM layers and 256 hidden units, while the student has one layer and 128 units. Both models are trained for 100 epochs with an initial learning rate of $1\times 10^{-3}$, which is halved when the validation loss does not improve for three consecutive epochs; training is stopped if no improvement is seen for ten epochs. We use Adam and apply gradient clipping with a maximum L2 norm of 5 and a batch size of 16.
@@ 5 Acknowledgements
This work is supported by the National Research Foundation of Korea (NRF) grant (No. RS-2024-00337945) and STEAM research grant (No. RS-2024-00464269) funded by the Ministry of Science and ICT of Korea government (MSIT).

### v2

@@ 1 Introduction
While deep neural networks have demonstrated strong performance in speech enhancement [4, 10, 16], their high computational demands pose significant challenges for practical applications, particularly in resource-constrained settings like on-device AI, where hardware limitations, power consumption, and monetary costs are critical concerns. Consequently, model compression has become a crucial research direction for developing computationally efficient speech enhancement models. Several compression strategies have been explored, including pruning [7], which removes redundant weights; quantization [3], which reduces the numerical precision of model parameters; and Knowledge Distillation (KD) [9, 22, 18]. This paper focuses on KD, a teacher-student paradigm in which a compact student model is trained to mimic a larger teacher model, thereby inheriting the teacher’s knowledge.
@@
KD has been actively explored in the speech domain, such as speaker verification [19, 24], speech separation [2, 27], and speech enhancement [8, 21, 25, 20, 26]. A prominent line of work is response-based KD, which trains a student to mimic the teacher’s output [19, 2, 27, 8, 26]. This output can be diverse, ranging from logits or posterior probabilities to time-frequency output. Some approaches [24, 21, 25] combine response-based and feature-based KD to align intermediate representations, leveraging complementary information.
@@
In speech enhancement, distilling time–frequency output has become a prominent response-based KD strategy because this approach can effectively capture the rich time-frequency structure of speech. This approach has given rise to notable frequency-dependent methods, such as Subband-KD [8] and Dynamic Frequency-Adaptive Knowledge Distillation (DFKD) [26]. DFKD addresses the limitation of Subband-KD arising from static band partitioning by introducing a frequency adapter that dynamically partitions the spectrogram into low- and high-frequency regions and optimizes each region with distinct losses. This dynamic frequency-dependent approach allows DFKD to significantly outperform Subband-KD and ABC-KD [25], a hybrid of response- and feature-based distillation. Although DFKD has shown strong performance, it remains limited by relying on the entire teacher output, which may contain low‑quality time–frequency regions.
@@
KD generally assumes that the teacher provides reliable supervision. However, this premise overlooks two critical issues in practice. The first limitation concerns the student’s existing proficiency. Distilling knowledge in regions where the student has already converged yields minimal benefit. The second challenge stems from the inherent imperfections of the teacher model, which may result from overfitting to the training dataset. Therefore, forcing the student to learn from flawed regions in the teacher’s output is detrimental, as it leads the student to replicate the teacher’s weakness instead of its strengths. As a result, indiscriminate distillation not only diverts the student’s focus from the regions where the teacher’s guidance is most needed but also, more importantly, can cause the student to imitate the teacher’s erroneous output, leading to a degradation in performance. Therefore, it is crucial to adopt a selective strategy that allows the student to learn only from beneficial supervision. Although several adaptive strategies have been proposed, such as importance-weighted KD [15, 6] and uncertainty-aware KD [12, 11], they still presuppose that the teacher is uniformly superior to the student for a given input. Consequently, they inevitably transfer local artifacts even under reduced weights, neglecting that the student may locally outperform the teacher.
@@
Fig. 1: Overview of DISPatch framework. DISPatch framework separates spectrograms into patches and selectively applies KD loss to those exhibiting a high Knowledge Gap, where the teacher’s error is low (L) and the student’s is high (H). Patch size: $2C\times N\times 1$ .
@@
Recent work on large language models has demonstrated that their accelerated convergence and improved performance can be achieved through selective training strategies [17]. This approach involves training the model exclusively on useful tokens, identified by a reference model trained on a high-quality dataset. Inspired by this principle, we propose Distilling Selective Patches (DISPatch). This strategy distills only the most useful spectrogram patches, leveraging KD’s selectivity and mitigating the limitations discussed above. The overall architecture of DISPatch is illustrated in Fig.1. We introduce a Knowledge Gap Score (KGS) to enable a more targeted transfer of the teacher’s knowledge by identifying and distilling only the most valuable patches, rather than imposing the KD loss across the entire spectrogram. DISPatch explicitly ignores regions where the teacher is unreliable or where the student is already proficient. Moreover, recognizing that speech characteristics vary across frequencies, we introduce Multi-Scale Selective Patches (MSSP). This is a frequency-dependent strategy that adapts the distinct patch sizes to different frequency regions for more tailored distillation. We show consistent gains by applying DISPatch to response-based KD, and further improvements when integrating DISPatch and MSSP into DFKD.
@@ 2.1 Not All Spectrogram Regions Are Equal
Fig. 2: Analysis of patch-level teacher-student misalignment. (a) Confusion matrix of patch classification. (b) Scatter plot comparing patch-wise errors between the teacher and the student.
@@
Contrary to the common assumption that the teacher consistently provides superior supervision, our patch-level analysis of a single speech sample reveals misalignment between the teacher and student. Fig.2 (a) visualizes the patch classification mismatch between the two models. Although the student has stabilized in many regions (Converged or Converging), a significant subset exists where the teacher remains unstable (Exploding or Challenging) despite the student’s proficiency (highlighted in the red box). Furthermore, Fig.2 (b) confirms that at the $J$-th point, the teacher’s error exceeds the student’s in 46.8% of patches. This empirical evidence exposes two critical risks of indiscriminate distillation: the transfer of harmful artifacts by forcing the student to mimic the teacher’s flawed regions, and the diversion of the student’s focus to regions where the student already excels, rather than where distillation is truly needed.
@@ 2.2 DISPatch: Distilling Selective Patches
2.3 MSSP: Multi-Scale Selective Patches
@@ 3.2 Implementation Details
We evaluate DISPatch using ConvTasNet (Deep w/ PReLU) [14] and DCCRN-CL [10]. For the implementation of ConvTasNet (Deep w/ PReLU), the teacher (9.71M) and student (1.75M) are configured with hyperparameters {N, L, B, H, Sc, P, X, R} set to {512, 16, 128, 512, 128, 3, 8, 3} and {128, 40, 128, 256, 128, 3, 7, 2}, respectively. The DCCRN-CL teacher (9.22M) and student (2.24M) are configured with encoder channels of {64, 128, 256, 256, 512, 512} and {32, 64, 128, 128, 256, 256}, respectively. Additionally, the teacher is configured with two complex LSTM layers and 256 hidden units, while the student has one layer and 128 units. Both models are trained for 100 epochs with an initial learning rate of $1\times 10^{-3}$, which is halved when the validation loss does not improve for three consecutive epochs; training is stopped if no improvement is seen for ten epochs. We use Adam and apply gradient clipping with a maximum L2 norm of 5 and a batch size of 16.
@@ 5 Acknowledgements
This work is supported by the National Research Foundation of Korea (NRF) grant (No. RS-2024-00337945) and STEAM research grant (No. RS-2024-00464269) funded by the Ministry of Science and ICT of Korea government (MSIT), the BK21 FOUR program through the NRF grant funded by the Ministry of Education of Korea government (MOE).